OpenStudy (anonymous):

Is this correct ? Factor: x2 - 8x + 16 answer is . . . (x - 8)(x + 8)

OpenStudy (anonymous):

false

OpenStudy (anonymous):

its not

OpenStudy (hoblos):

x2 - 8x + 16 =(x-4)^2 so no it is not correct

OpenStudy (anonymous):

-4 and -4 are the factors of 16 which add up to -8

OpenStudy (anonymous):

(x-8)(x+8) = '' x2 +64 ''

OpenStudy (hoblos):

@Saleh101 the difference of two squares formula says (a-b)(a+b) = a^2 - b^2 minus NOt plus !!

OpenStudy (anonymous):

yep my bad, its funny i always end up with such mistakes. thanks

OpenStudy (hoblos):

no problem everybody makes mistakes.. just try to be careful ;)